# A physics-compliant diagonal representation for wireless channels parametrized by beyond-diagonal reconfigurable intelligent surfaces

### v1

@@ III-B Numerical full-wave simulations
Whereas the previous subsection covered theoretical ideal load circuits that can be analyzed analytically, we now turn our attention to load circuits that can be realized in practise. These can generally not avoid the effects of loss, delay and dispersion, and in most cases, an analytical expression for $\mathbf{S}^{\mathrm{LC}}$ will not exist. Nonetheless, it is reasonable to assume that for practical studies of realistic BD-RIS designs a detailed layout of the circuit terminating the RIS elements is available. Indeed, such a layout will be an inevitable pre-requisite for any experimentally realized prototype fabrication. Consequently, this circuit can be simulated numerically in full-wave solvers. In fact, a single full-wave simulation is sufficient to obtain $\mathbf{S}^{\mathrm{LC}}$ if one simply places lumped ports at the locations of the tunable lumped elements. Incidentally, this is exactly the procedure that was already implemented in Ref. [28] to extract the analytically intractable scattering matrix of a highly complex on-chip radio environment parametrized by a D-RIS. As mentioned earlier, a radio environment parametrized by a D-RIS is conceptually exactly the same problem as a load circuit parametrized by tunable individual loads. To summarize, for any arbitrarily complex load circuit design, a single full-wave simulation can yield $\mathbf{S}^{\mathrm{LC}}$.
@@ III-C Experimental estimation
In this subsection, we discuss the extent to which $\mathbf{S}^{\mathrm{LC}}$ could be estimated experimentally. A pivotal requirement for such an experimental estimation would be the possibility to input and output waves through the load circuit ports connected to the RIS elements (i.e., ports whose indices are in the set $\bar{\mathcal{S}}$); this could be possible, for instance, before connecting the RIS and the load circuit. The considered problem is analogous to the problem of physics-compliant end-to-end channel estimation with a D-RIS that was tackled in Refs. [19, 30]. Indeed, the load circuit is a multi-port network characterized by $\mathbf{S}^{\mathrm{LC}}$ that is terminated by individual tunable loads. Based on experimental measurements at the ports of the load circuit from set $\bar{\mathcal{S}}$ for different terminations of the ports from set $\mathcal{C}$, it is provably impossible to unambiguously determine $\mathbf{S}^{\mathrm{LC}}$ experimentally because at least sign ambiguities on off-diagonal terms associated with the ports from set $\mathcal{C}$ will remain [30, 16]. If not at least three distinct and characterized individual load states are available at the ports from $\mathcal{C}$, additional ambiguities about $\mathbf{S}^{\mathrm{LC}}$ ensue. However, operationally, these ambiguities are not problematic because as long as $\mathbf{S}^{\mathrm{L}}$ is correctly predicted we have everything we need to terminate the ports of $\mathbf{S}^{\mathrm{RE}}$ from set $\mathcal{S}$ and thus determine $\tilde{\mathbf{S}}$ and $\mathbf{H}$. Indeed, Ref. [19] accurately predicted wireless end-to-end channels in an unknown complex medium as a function of the binary configuration of a D-RIS prototype with unknown characteristics. The infinite set of possible parameter values should even facilitates the estimation of a plausible $\mathbf{S}^{\mathrm{LC}}$ (and $\mathbf{S}^{\mathrm{IL}}$ if it is unknown) that yields an accurate prediction for $\mathbf{S}^{\mathrm{L}}$.
III-D Do we need to know $\mathbf{S}^{\mathrm{LC}}$ explicitly?
@@
This section was dedicated to discussing various routes to determining $\mathbf{S}^{\mathrm{LC}}$, a quantity that is at the core of the physics-compliant diagonal representation of BD-RIS parametrized channels put forth in the present paper. Although this section is pedagogically important to substantiate the theory developed in the previous section, arguably, in practise it is not even necessary to explicitly know $\mathbf{S}^{\mathrm{LC}}$.
@@
Indeed, in practise, the most compact modeling approach would directly map the configuration of the individual tunable loads to the end-to-end wireless channels. The primary entities are the antennas and the tunable lumped elements. Meanwhile, the RIS elements are merely a detail of how these primary entities are coupled to each other. As long as we know $\mathbf{S}^{\mathrm{casc}}$, operationally there is no need to break it down into $\mathbf{S}^{\mathrm{RE}}$ and $\mathbf{S}^{\mathrm{LC}}$. Therefore, in the next section, we propose a scheme for end-to-end channel estimation and optimization that only estimates $\mathbf{S}^{\mathrm{casc}}$ in order to successfully predict $\mathbf{H}$ as a function of the tunable individual loads.
@@ IV Directly applying D-RIS channel estimation and optimization algorithms to BD-RIS
This section is dedicated to demonstrating the feasibility of directly applying channel estimation and optimization algorithms originally developed for D-RIS to BD-RIS-parametrized channels, based on an experimentally grounded case study. We first describe the considered scenario, then channel estimation, and finally channel optimization.
@@ IV-A Considered scenario grounded in an experiment
To ground the considered scenario in an experimental reality involving unknown rich scattering, we have placed eight commercial antennas (AEACBK081014-S698, designed for operation between 698 MHz and 2.69 GHz for GSM, SPRS and 4G (LTE)) inside a reverberation chamber (1.75 m $\times$ 1.50 m $\times$ 2.00 m) and measured the corresponding $8\times 8$ scattering matrix with an eight-port vector network analyzer (Keysight M9005A) at $f=800$ MHz. The obtained scattering matrix is used as $\mathbf{S}^{\mathbf{RE}}$ in this case study. We consider a SISO link involving one transmitting and one receiving antenna, and the remaining six antennas are the RIS elements. The annotations on the photographic image in Fig. 3 highlight this setup. As sketched in the right part of Fig. 3, the six RIS elements are split into three groups of size two that are considered to be connected to ideal $\pi$ networks, each of which is parametrized by three impedances. We assume that each impedance represents a PIN diode with two possible states whose characteristics are chosen according to those documented for a commercially available PIN diode (MADP-000907-14020x [46]). Specifically, the two possible load impedances are $\eta^{\mathrm{ON}}=5.2\leavevmode\nobreak\ \Omega$ and $\eta^{\mathrm{OFF}}=(\jmath\omega C)^{-1}=-7.96\jmath\times 10^{3}\leavevmode\nobreak\ \Omega$, where $\jmath=\sqrt{-1}$, $\omega=2\pi f$ and $C=25\ \mathrm{fF}$. The corresponding reflection coefficients are $r_{1}=(\eta^{\mathrm{ON}}-Z_{0})/(\eta^{\mathrm{ON}}+Z_{0})=-0.81$ and $r_{2}=(\eta^{\mathrm{OFF}}-Z_{0})/(\eta^{\mathrm{OFF}}+Z_{0})=1.00-0.01\jmath$. To recapitulate key parameters: $N_{\mathrm{T}}=1$, $N_{\mathrm{R}}=1$, $N_{\mathrm{A}}=N_{\mathrm{T}}+N_{\mathrm{R}}=2$, $N_{\mathrm{S}}=6$ and $N_{\mathrm{C}}=9$.
@@
Our experimental measurements hence fix the entries of $\mathbf{S}^{\mathrm{RE}}$, Eq. (13) fixes the entries of the three diagonal blocks of the block-diagonal matrix $\mathbf{S}^{\mathrm{LC}}$, and the two possible values that the diagonal entries of the diagonal matrix $\mathbf{S}^{\mathrm{IL}}$ can take are $r_{1}$ and $r_{2}$ derived from the PIN diode’s data sheet. We can consequently evaluate $\mathbf{S}^{\mathrm{L}}$ using Eq. (1) or via the known admittance matrix of an ideal $\pi$ network. We can also evaluate $\mathbf{S}^{\mathrm{casc}}$ using Eq. (8). Finally, we can determine $\tilde{\mathbf{S}}$ (and thus $\mathbf{H}$) using Eq. (2) based on $\mathbf{S}^{\mathrm{RE}}$ and $\mathbf{S}^{\mathrm{L}}$ or using Eq. (10) based on $\mathbf{S}^{\mathrm{casc}}$ and $\mathbf{S}^{\mathrm{IL}}$. For the considered SISO problem, $\mathbf{H}$ is a scalar (the off-diagonal entry of $\tilde{\mathbf{S}}$) such that we denote it by $h$ in the following. For a given binary configuration $\mathbf{b}\in\mathbb{C}^{N_{\mathrm{C}}\times 1}$ of the nine tunable load impedances, we can hence identify the corresponding $\mathbf{S}^{\mathrm{IL}}$ and ultimately the corresponding $h$. We verified that both the conventional and the alternative route to evaluating $h$ yields exactly the same value for multiple random choices of $\mathbf{b}$. Altogether, to the best of our knowledge, this constitutes the first study of a BD-RIS-parametrized wireless channel for which the utilized parameters are grounded in an experimental reality.
@@
To perform physics-compliant end-to-end channel estimation and optimization, we simulate a scenario in which we can configure the reconfigurable impedance network to any desired binary configuration $\mathbf{b}$ and measure the corresponding wireless channel $h(\mathbf{b})$. Of course, we obtain $h(\mathbf{b})$ as detailed in the previous paragraph, but once the pairs $\{\mathbf{b},h(\mathbf{b})\}$ are generated, we assume that we only know the pairs $\{\mathbf{b},h(\mathbf{b})\}$ and do not have any knowledge of $r_{1}$, $r_{2}$, $\mathbf{S}^{\mathrm{LC}}$ or $\mathbf{S}^{\mathrm{RE}}$. We thereby simulate a true end-to-end channel estimation and optimization problem without any a priori knowledge of the characteristics of the radio environment, the load circuit or the tunable lumped elements.
IV-B Channel estimation
@@
The literature contains a few theoretical studies on channel estimation in scenarios involving BD-RIS [47, 48]. However, these are based on simplified cascaded channel models. Here, we tackle the problem of physics-compliant end-to-end channel estimation of BD-RIS parametrized channels in a scenario that is grounded in experimental data. We highlight that thanks to the diagonal representation of the BD-RIS-parametrized channel we only need to estimate some version of $\mathbf{S}^{\mathrm{casc}}$ and $\mathbf{S}^{\mathrm{IL}}$ but not the breakdown of $\mathbf{S}^{\mathrm{casc}}$ into $\mathbf{S}^{\mathrm{RE}}$ and $\mathbf{S}^{\mathrm{LC}}$. (The words “some version” in the previous sentence hint at the fact that ambiguities are inevitable and not problematic, as discussed below.)
@@ IV-B1 Problem statement
We hence face a parameter estimation problem. The most compact approach is based on the diagonal representation of the BD-RIS-parametrized channel. It consists in determining a set of parameters comprising the entries of $\mathbf{S}^{\mathrm{casc}}$ as well as the two possible values $r_{1}$ and $r_{2}$ that each diagonal entry of $\mathbf{S}^{\mathrm{IL}}$ can take such that we correctly predict $h$ using Eq. (10) and Eq. (3) for any given $\mathbf{b}$.
@@
    $\begin{split}N_{\mathrm{params}}&=2\left(2+\frac{(N_{\mathrm{A}}+N_{\mathrm{C}})(N_{\mathrm{A}}+N_{\mathrm{C}}+1)}{2}\right)\\ &=4+(N_{\mathrm{A}}+N_{\mathrm{C}})(N_{\mathrm{A}}+N_{\mathrm{C}}+1)\\ &=136,\end{split}$ (15)
    where the first term inside the brackets corresponds to the two possible reflection coefficients $r_{1}$ and $r_{2}$ of the individually tunable lumped elements, the second term inside the brackets corresponds to the entries of $\mathbf{S}^{\mathrm{casc}}$ which is symmetric due to reciprocity, and the factor two in front of the brackets takes into account that the variables are complex-valued. Note that the number of parameters to be estimated neither depends on $N_{\mathrm{S}}$ nor on the complexity of the radio environment nor on the complexity of the load circuit.
@@ IV-B2 Algorithm
Our physics-compliant end-to-end channel estimation algorithm closely follows that proposed in Sec. IV of Ref. [30]. However, our present problem differs from that considered in Ref. [30] in that we only measure $h$ instead of $\tilde{\mathbf{S}}$ and in that we do not know the values of $r_{1}$ and $r_{2}$. Hence, we seek to estimate the entries of $\mathbf{S}^{\mathrm{casc}}$ (except for the diagonal ones in the block $\mathcal{AA}$) as well as $r_{1}$ and $r_{2}-r_{1}$ simultaneously. The earlier gradient-descent proposal from Ref. [19] did not use an advantageous two-step procedure but it did also treat a problem with unknown load characteristics. Our algorithmic implementation consists simply in optimizing the sought-after parameters via gradient descent such that a cost function quantifying the difference between the predicted and ground truth values of the difference of $h$ between successive configurations of the individual tunable loads is minimized. Additional implementation details can be found in Ref. [30].

### v2

@@ III-B Numerical full-wave simulations
Whereas the previous subsection covered theoretical ideal load circuits that can be analyzed analytically, we now turn our attention to load circuits that can be realized in practice. These can generally not avoid the effects of loss, delay and dispersion, and in most cases, an analytical expression for $\mathbf{S}^{\mathrm{SLC}}$ will not exist. Nonetheless, it is reasonable to assume that for practical studies of realistic BD-RIS designs a detailed layout of the circuit terminating the RIS elements is available. Indeed, such a layout will be an inevitable pre-requisite for any experimentally realized prototype fabrication. Consequently, this circuit can be simulated numerically in full-wave solvers. In fact, a single full-wave simulation is sufficient to obtain $\mathbf{S}^{\mathrm{SLC}}$ if one simply places lumped ports at the locations of the tunable lumped elements. Incidentally, this is exactly the procedure that was already implemented in Ref. [28] to extract the analytically intractable scattering matrix of a highly complex on-chip RE parametrized by a D-RIS. As mentioned earlier, an RE parametrized by a D-RIS is conceptually exactly the same problem as a load circuit parametrized by tunable individual loads. To summarize, for any arbitrarily complex load circuit design, a single full-wave simulation can yield $\mathbf{S}^{\mathrm{SLC}}$.
@@ III-C Experimental estimation
In this subsection, we discuss the extent to which $\mathbf{S}^{\mathrm{SLC}}$ could be estimated experimentally. A pivotal requirement for such an experimental estimation would be the possibility to input and output waves through the load circuit ports connected to the RIS elements (i.e., ports of SLC whose indices are in the set $\bar{\mathcal{S}}$); this could be possible, for instance, before connecting the RIS and the load circuit. Then, the considered problem is analogous to the problem of physics-compliant end-to-end channel estimation with a D-RIS that was tackled in Refs. [19, 30]. Indeed, the load circuit is a multi-port network characterized by $\mathbf{S}^{\mathrm{SLC}}$ that is terminated by individual tunable loads. Based on experimental measurements at the ports of the load circuit from set $\bar{\mathcal{S}}$ for different terminations of the ports from set $\mathcal{C}$, it is provably impossible to unambiguously determine $\mathbf{S}^{\mathrm{SLC}}$ experimentally because at least sign ambiguities on off-diagonal terms associated with the ports from set $\mathcal{C}$ will remain [30, 16]. If not at least three distinct and characterized individual load states are available at the ports from $\mathcal{C}$, additional ambiguities about $\mathbf{S}^{\mathrm{SLC}}$ ensue. However, operationally, these ambiguities are not problematic because as long as $\mathbf{S}^{\mathrm{L}}$ is correctly predicted we have everything we need to evaluate $\mathbf{H}$ based on Eq. (2) by cascade-loading $\mathbf{S}^{\mathrm{RE}}$ at its ports comprised in $\mathcal{S}$ with $\mathbf{S}^{\mathrm{L}}$. Indeed, Ref. [19] accurately predicted wireless end-to-end channels in an unknown complex medium as a function of the binary configuration of a D-RIS prototype with unknown characteristics. The infinite number of satisfactory sets of parameter values should even facilitates the estimation of a plausible $\mathbf{S}^{\mathrm{SLC}}$ (and $\mathbf{S}^{\mathrm{IL}}$ if it is unknown) that yields an accurate prediction for $\mathbf{S}^{\mathrm{L}}$.
III-D Do we need to know $\mathbf{S}^{\mathrm{SLC}}$ explicitly?
@@
This section was dedicated to discussing various routes to determining $\mathbf{S}^{\mathrm{SLC}}$, a quantity that is at the core of the physics-compliant diagonal representation of BD-RIS parametrized channels put forth in the present paper. Although this section is pedagogically important to substantiate the theory developed in the previous section, arguably, in practice it is not even necessary to explicitly know $\mathbf{S}^{\mathrm{SLC}}$.
@@
Indeed, in practice, the most compact modeling approach is our diagonal representation from Eq. (10) which directly maps the configuration of the individual tunable loads to the end-to-end wireless channels without separating K into RE and SLC. As long as we know $\mathbf{S}^{\mathrm{K}}$, operationally there is no need to break it down into $\mathbf{S}^{\mathrm{RE}}$ and $\mathbf{S}^{\mathrm{SLC}}$. The primary ports in a BD-RIS chain cascade (see Fig. 1) are those associated with the antennas and the tunable lumped elements. Meanwhile, the RIS elements (and the associated ports) are merely a detail of how these primary entities are coupled to each other. Therefore, in the next section, we propose a scheme for end-to-end channel estimation and optimization that only estimates $\mathbf{S}^{\mathrm{K}}$ and $\mathbf{S}^{\mathrm{IL}}$ in order to successfully predict $\mathbf{H}$ as a function of the BD-RIS configuration.
@@ IV Directly applying D-RIS channel estimation and optimization algorithms to BD-RIS
The most significant implication of our diagonal representation of BD-RIS parametrized channels for signal processing and wireless communications is that it implies that physics-compliant D-RIS algorithms can be directly applied in BD-RIS scenarios. This section is dedicated to demonstrating the feasibility of directly applying physics-compliant end-to-end channel estimation and optimization algorithms originally developed for D-RIS to BD-RIS-parametrized channels, based on an experimentally grounded case study. We first describe the considered scenario, then channel estimation, and finally channel optimization.
@@ IV-A Considered scenario grounded in an experiment
To ground the considered scenario in an experimental reality involving unknown rich scattering, we have placed eight commercial antennas (AEACBK081014-S698, designed for operation between 698 MHz and 2.69 GHz for GSM, SPRS and 4G (LTE)) inside a reverberation chamber (1.75 m $\times$ 1.50 m $\times$ 2.00 m) and measured the corresponding $8\times 8$ scattering matrix with an eight-port vector network analyzer (Keysight M9005A) at $f=800$ MHz. The obtained scattering matrix is used as $\mathbf{S}^{\mathrm{RE}}$ in this case study. We consider a SISO link involving one transmitting and one receiving antenna, and the remaining six antennas are the RIS elements. The annotations on the photographic image in Fig. 3 highlight this setup. As sketched in the right part of Fig. 3, the six RIS elements are split into three groups of size two that are considered to be connected to ideal $\pi$ networks, each of which is parametrized by three impedances. We assume that each impedance represents a PIN diode with two possible states whose characteristics are chosen according to those documented for a commercially available PIN diode (MADP-000907-14020x [50]). Specifically, the two possible load impedances are $\eta^{\mathrm{ON}}=5.2\leavevmode\nobreak\ \Omega$ and $\eta^{\mathrm{OFF}}=(\jmath\omega C)^{-1}=-7.96\jmath\times 10^{3}\leavevmode\nobreak\ \Omega$, where $\jmath=\sqrt{-1}$, $\omega=2\pi f$ and $C=25\ \mathrm{fF}$. The corresponding reflection coefficients are $r_{\mathrm{A}}=(\eta^{\mathrm{ON}}-Z_{0})/(\eta^{\mathrm{ON}}+Z_{0})=-0.81$ and $r_{\mathrm{B}}=(\eta^{\mathrm{OFF}}-Z_{0})/(\eta^{\mathrm{OFF}}+Z_{0})=0.9999-0.0126\jmath$. To recapitulate key parameters: $N_{\mathrm{T}}=1$, $N_{\mathrm{R}}=1$, $N_{\mathrm{A}}=N_{\mathrm{T}}+N_{\mathrm{R}}=2$, $N_{\mathrm{S}}=6$ and $N_{\mathrm{C}}=9$.
Our experimental measurements hence fix the entries of $\mathbf{S}^{\mathrm{RE}}$, Eq. (19) fixes the entries of the three diagonal blocks of the block-diagonal matrix $\mathbf{S}^{\mathrm{SLC}}$, and the PIN diode’s data sheet fixes the two possible values that the diagonal entries of the diagonal matrix $\mathbf{S}^{\mathrm{IL}}$ can take. Specifically, for a given binary configuration $\mathbf{b}\in\mathbb{B}^{N_{\mathrm{C}}\times 1}$ of the $N_{\mathrm{C}}=9$ tunable load impedances, we can hence identify the corresponding $\mathbf{S}^{\mathrm{IL}}$ via
$\mathbf{S}^{\mathrm{IL}}(\mathbf{b})=\mathrm{diag}\left(r_{\mathrm{A}}\mathbf{1}_{N_{\mathrm{C}}}+(r_{\mathrm{B}}-r_{\mathrm{A}})\mathbf{b}\right).$ (20)
@@
We can evaluate $\mathbf{S}^{\mathrm{L}}(\mathbf{b})$ by inserting $\mathbf{S}^{\mathrm{IL}}(\mathbf{b})$ (obtained with Eq. (20)) into Eq. (5) or via the known admittance matrix of an ideal $\pi$ network. Moreover, we can evaluate $\mathbf{S}^{\mathrm{K}}$ using Eq. (7). Finally, we can determine $\mathbf{H}(\mathbf{b})$ via the conventional route using Eq. (2) based on $\mathbf{S}^{\mathrm{RE}}$ and $\mathbf{S}^{\mathrm{L}}(\mathbf{b})$ or via our alternative route using Eq. (10) based on $\mathbf{S}^{\mathrm{K}}$ and $\mathbf{S}^{\mathrm{IL}}(\mathbf{b})$. We verified that both routes to evaluating $\mathbf{H}(\mathbf{b})$ yield exactly the same result for multiple random choices of $\mathbf{b}$. For the considered SISO problem, $\mathbf{H}(\mathbf{b})$ is a scalar such that we denote it by $h(\mathbf{b})$ in the following. Altogether, to the best of our knowledge, this constitutes the first study of a BD-RIS-parametrized wireless channel for which the utilized parameters are grounded in an experimental reality.
@@
To perform physics-compliant end-to-end channel estimation and optimization, we simulate a scenario in which we can configure the reconfigurable impedance network to any desired binary configuration $\mathbf{b}$ and measure the corresponding wireless channel $h(\mathbf{b})$. Of course, we obtain $h(\mathbf{b})$ as detailed in the previous paragraph, but once the pairs $\{\mathbf{b},h(\mathbf{b})\}$ are generated, we assume that we only know the pairs $\{\mathbf{b},h(\mathbf{b})\}$ and do not have any knowledge of $\mathbf{S}^{\mathrm{RE}}$, $\mathbf{S}^{\mathrm{SLC}}$, $r_{\mathrm{A}}$, or $r_{\mathrm{B}}$. We thereby simulate a true end-to-end channel estimation and optimization problem without any a priori knowledge of the characteristics of the radio environment, the load circuit or the tunable lumped elements.
IV-B End-to-end channel estimation
@@
The literature contains a few theoretical studies on channel estimation in scenarios involving BD-RIS [51, 52]. However, these are based on simplified cascaded channel models. Here, we tackle the problem of physics-compliant end-to-end channel estimation of BD-RIS parametrized channels in a scenario that is grounded in experimental data. We highlight that thanks to the diagonal representation of the BD-RIS-parametrized channel we only need to estimate some version of $\mathbf{S}^{\mathrm{K}}$ and $\mathbf{S}^{\mathrm{IL}}$ but not the breakdown of $\mathbf{S}^{\mathrm{K}}$ into $\mathbf{S}^{\mathrm{RE}}$ and $\mathbf{S}^{\mathrm{SLC}}$. (The words “some version” in the previous sentence hint at the fact that ambiguities are inevitable and not problematic, as discussed below.)
@@ IV-B1 Problem statement
We hence face a parameter estimation problem. The most compact approach is based on the diagonal representation of the BD-RIS-parametrized channel in Eq. (10). It consists in determining the parameters necessary to predict $h$ using Eq. (10) for any given $\mathbf{b}$. These parameters are the blocks $\mathcal{RT}$, $\mathcal{AC}$ and $\mathcal{CC}$ of $\mathbf{S}^{\mathrm{K}}$ as well as the two possible values $r_{\mathrm{A}}$ and $r_{\mathrm{B}}$ that each diagonal entry of $\mathbf{S}^{\mathrm{IL}}$ can.
@@
    $\begin{split}N_{\mathrm{params}}&=2\left(2+\left[\frac{(N_{\mathrm{A}}+N_{\mathrm{C}})(N_{\mathrm{A}}+N_{\mathrm{C}}+1)}{2}-2\right]\right)\\ &=4+\left[(N_{\mathrm{A}}+N_{\mathrm{C}})(N_{\mathrm{A}}+N_{\mathrm{C}}+1)-4\right]\\ &=132,\end{split}$ (22)
    where the first term inside the brackets corresponds to the two possible reflection coefficients $r_{\mathrm{A}}$ and $r_{\mathrm{B}}$ of the individually tunable lumped elements, the second term inside the brackets corresponds to the number of unique entries in the blocks $\mathcal{RT}$, $\mathcal{AC}$ and $\mathcal{CC}$ of $\mathbf{S}^{\mathrm{K}}$, and the factor two in front of the brackets takes into account that the variables are complex-valued. Note that the number of parameters to be estimated neither depends on $N_{\mathrm{S}}$ nor on the complexity of the radio environment nor on the complexity of the load circuit.
@@ IV-B2 Algorithm
Our physics-compliant end-to-end channel estimation algorithm closely follows that proposed in Sec. IV of Ref. [30]. However, our present problem differs from that considered in Ref. [30] in that we only measure $h$ instead of $\tilde{\mathbf{S}}$ and in that we do not know the values of $r_{\mathrm{A}}$ and $r_{\mathrm{B}}$. Hence, we seek to estimate the entries of $\mathbf{S}^{\mathrm{K}}$ (except for the diagonal ones in the block $\mathcal{AA}$) as well as $r_{\mathrm{A}}$ and $r_{\mathrm{B}}-r_{\mathrm{A}}$ simultaneously. The earlier gradient-descent proposal from Ref. [19] did not use an advantageous two-step procedure but it did also treat a problem with unknown load characteristics. Our algorithmic implementation consists simply in optimizing the sought-after parameters via gradient descent such that a cost function quantifying the difference between the predicted and ground-truth values of the difference of $h$ between successive configurations of the individual tunable loads is minimized. Additional implementation details can be found in Ref. [30].
The algorithm is not specific to a SISO scenario; in fact, the performance of the algorithm improves the more antennas on transmitting and receiving sides are involved, as seen in Fig. 2 in Ref. [19]. We merely chose a SISO scenario in this case study because of the limited number of antennas in our experiment (see Fig. 3A), and because of the ease of defining the cost function in terms of the RSSI in the SISO case.